# Distributed implementation of tensor-product selected configuration interaction

### v1

@@ 2.4 Strategies for Efficient Parallel Execution in TPSCI
As described above, we distribute the CI vector across nodes to overcome single-node memory limitations, enabling computations to fully utilize the total available memory. However, this distributed storage introduces new challenges: a computation that could previously be completed in a single step now requires $N$ steps, with each step requiring data retrieval from a target process [as shown in Eq. (3)]. As the number of nodes increases, the overall MPI communication volume seemingly scales as $N_{\text{SCI}}^{2}$, making communication latency a potential bottleneck for computational efficiency. To address this issue, we have developed a series of strategies to reduce communication overhead and improve parallel execution.
@@ 2.4.1 Avoiding Unnecessary MPI Data Transfers
In our previous analysis, we mentioned that each local process $p$ needs to fetch data from other processes. However, not all data from the target process is required. Specifically, if all $\alpha$ bit strings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{u})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{u}$.
In our implementation, such information is precomputed and stored, allowing each process to identify and skip unnecessary data transfers, thereby significantly improving communication efficiency. As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot y$. It can be easily shown that the computational cost and the data communication volume on each process share the same scaling. Consequently, the overall MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$ rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TPSCI on supercomputers.
@@ 2.4.2 Minimizing Long-Distance Communication in MPI
Each node in the supercomputer is assigned a unique node identity number (ID) ranging from 00 to $N-1$. When suitably ordered, messages sent between nodes with closer IDs will on average require fewer hops compared to those with larger ID differences.
@@
Based on this, we sort the $\alpha$ bit strings according to their excitation levels relative to the Hartree-Fock (HF) $\alpha$ bit string, and then assign them to nodes in increasing order. Hence, lower-excitation $\alpha$ bit strings are assigned to lower-numbered nodes, whereas higher-excitation strings are assigned to higher-numbered nodes.
As a result, communication between distant nodes is significantly reduced because the excitation differences between their assigned $\alpha$ bit strings are more likely to be all greater than 2, making communication unnecessary. This strategy is crucial for reducing MPI communication congestion, a well-known challenge in large-scale high-performance computing, by ensuring that each node primarily fetches data from nearby nodes, thereby improving parallel efficiency.
2.4.3 Achieving Computational Load Balance in MPI
In MPI-based parallel computing, maintaining computational load balance is crucial because the slowest process determines the overall execution time. The challenge is how to estimate each process’s workload in advance. As shown in Algorithm 1, the most time-consuming computation is the evaluation of the [1,1] term. Therefore, the number of single excitations associated with each $\alpha$ bit string can be used as a rough estimate of the computational workload for its corresponding segment. Consequently, the total computational workload of a process can be estimated by summing the single excitation counts of all $\alpha$ bit strings assigned to it.
As described above, we assign the $\alpha$ bit strings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bit string. This ordering is crucial in our TPSCI program to minimize long-distance communication between nodes. However, balancing memory usage and computational workload simultaneously presents a challenge. Since the total number of $\alpha$ bit strings is fixed, the average number per node is also predetermined. Similarly, the total number of single excitation counts of all $\alpha$ bit strings is fixed, meaning that the average single excitation count per process is also predetermined.
@@ 2.4.3 Achieving Computational Load Balance in MPI
To quantify the balance between memory and computation, we define two expansion factors: the memory expansion factor, which is the ratio of the assigned $\alpha$ bit string count to the average, and the computation expansion factor, which is the ratio of the assigned single excitation counts to the average. Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can lead to severe computational imbalance, causing some nodes to finish significantly earlier and remain idle while waiting for others. Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may result in a highly uneven memory distribution, with some nodes exceeding their memory limits, forcing an increase in the total number of nodes and leading to inefficient resource usage.
@@
To address this issue, we adopt a balanced strategy where both expansion factors are kept moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive.
2.4.4 Overlapping Computation and Data Transfer to Reduce MPI Latency
@@
To mitigate this issue, one core is assigned to perform MPI data transfers, fetching the required data for the next step, while the remaining 47 cores perform parallel computation for the current step using OpenMP. Let $t_{\text{cal}}$ and $t_{\text{data}}$ represent the computation time and the data retrieval time of one step, respectively. If $t_{\text{data}}>t_{\text{cal}}$, we define the delay time to the wall time as $t_{\text{delay}}=t_{\text{data}}-t_{\text{cal}}$. Conversely, if $t_{\text{data}}\leq t_{\text{cal}}$, then $t_{\text{delay}}=0$, since MPI communication congestion during data transfer—if it occurs—does not prolong the wall time. The total delay time for a Davidson iteration, denoted as $T_{\text{delay}}$, is given by the sum of $t_{\text{delay}}$ over all $N$ steps.
@@ 3 Results
Although the TPSCI program in this work is designed for “QC+HPC”, this study is an independent effort and does not yet incorporate QC-based bit string selection. Instead, we perform FCI calculations to provide widely applicable benchmark data. Consequently, the computational scaling follows $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}$, which is higher than the $N_{\mathrm{FCI}}\cdot N_{\mathrm{occ}}\cdot N_{\mathrm{vir}}$ scaling of the Handy-Knowles FCI algorithm. Our goal is not to compete with existing FCI methods but to showcase the effectiveness of the newly designed TPSCI algorithm and its high efficiency in parallel execution.
In actual calculations, we further introduced molecular symmetry to reduce both memory usage and computational cost. All HF reference states and their corresponding FCIDUMP files were generated using PySCF. 19 For double-precision TPSCI, convergence was defined as an energy difference of less than $1\times 10^{-7}$ Hartree between consecutive iterations. For single-precision TPSCI, due to fewer significant digits, a looser criterion of less than $1\times 10^{-6}$ Hartree was used. Notably, the final total energies obtained from both versions differ only in the sixth decimal place, indicating that the single-precision SCI calculation is accurate to $0.01$ milliHartree (mH), which is sufficient for chemical applications.
We perform experiments on supercomputer Fugaku, which is a CPU only architecture based on the A64FX processor. Each node has 48 cores, and only 32 GiB of memory, making processing and memory distribution essential. The code was compiled with the Fujitsu compilers using the -Kfast,parallel,openmp,SVE options.
@@ 3.1 N2 with the aug-cc-pVDZ Basis Set
Using $D_{2h}$ symmetry with frozen core orbitals, the N2 molecule with the aug-cc-pVDZ basis set has $1.47\times 10^{11}$ determinants. We first evaluate the computational time of a single Davidson iteration using different numbers of nodes. As shown in Table 1, the program’s wall time, defined as the wall time of the slowest process, remains within 116% of the average wall time across all processes, demonstrating good MPI load balancing.
@@
Table 1: Parallel performance of TPSCI for N2 computed with the aug-cc-pVDZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.

### v2

@@ 2.4 Strategies for Efficient Parallel Execution in TBSCI
As described above, we distribute the CI vector across nodes to overcome single-node memory limitations, enabling computations to fully utilize the total available memory. However, this distributed storage introduces new challenges: a computation that could previously be completed in a single step now requires $N$ steps, with each step requiring data retrieval from a target process [as shown in Eq. (3)]. As the number of nodes increases, the overall MPI communication volume seemingly scales as $N_{\text{SCI}}^{2}$, making communication latency a potential bottleneck for computational efficiency. To address this issue, we have developed a suite of strategies to reduce communication overhead and improve parallel execution.
@@ 2.4.1 Avoiding Unnecessary MPI Data Transfers
In our previous analysis, we mentioned that each local process $p$ needs to fetch data from other processes. However, not all data from the target process is required. Specifically, if all $\alpha$ bitstrings assigned to process $p$, namely $\{S^{\alpha}_{p,1},S^{\alpha}_{p,2},\dots,S^{\alpha}_{p,m_{p}}\}$, satisfy $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{x})>2$ for any $S^{\alpha}_{w}$ in this set, then process $p$ does not need the segment corresponding to $S^{\alpha}_{x}$.
In our implementation, such information is precomputed and stored, allowing each process to identify and skip unnecessary data transfers, thereby significantly improving communication efficiency. As a result, the MPI communication volume for process $p$ roughly scales as $m_{p}\cdot N_{\mathrm{occ}}^{2}\cdot N_{\mathrm{vir}}^{2}\cdot\sqrt{N_{\mathrm{SCI}}/N_{\mathrm{FCI}}}\cdot l_{\beta}$, matching the scaling of the local computational cost. Consequently, the total MPI communication volume across all nodes scales as $N_{\mathrm{SCI}}$, rather than $N_{\mathrm{SCI}}^{2}$, providing a solid foundation for the efficient parallel execution of TBSCI in large-scale calculations involving many nodes on supercomputers.
2.4.2 Exploiting Molecular Symmetry
Exploiting molecular symmetry is a well-established practice in quantum chemistry to enhance computational efficiency. Building upon the strategy of avoiding unnecessary MPI data transfers, leveraging molecular symmetry can further reduce the data transfer volume in the TBSCI framework. As shown in Algorithm 1, the segments satisfying $\text{DIFF}(S^{\alpha}_{w},S^{\alpha}_{x})=2$, which are typically much more than those satisfying $\text{DIFF}=1$, are used solely for computing the $[2,0]$ term. In this case, because the $\beta$ bitstrings must match exactly ($S^{\beta}_{v}\equiv S^{\beta}_{u}$), $S^{\alpha}_{w}$ and $S^{\alpha}_{x}$ must belong to the same symmetry sector, otherwise the corresponding Hamiltonian matrix element vanishes. Thus, molecular symmetry naturally restricts the possible communication targets and further reduces MPI communication. This strategy becomes particularly effective for molecules with high symmetry. For example, in systems with D2h symmetry, memory usage drops by $\sim$8×8$\times$8 ×, and computational cost and data transfer volume drop by $\sim$64×64$\times$64 ×.
However, in this case, the selected determinants are no longer constructed as tensor products of independently selected $\alpha$ and $\beta$ bitstrings, but are further restricted—on top of the original tensor-product construction—to those consistent with the target symmetry. As a result, the storage layout no longer strictly follows the structure shown in Figure 1, and the computation is no longer performed strictly according to Algorithm 1, as only symmetry-allowed combinations are evaluated. Nevertheless, for clarity, we retain the original forms of both Figure 1 and Algorithm 1.
@@ 2.4.3 Minimizing Long-Distance Communication in MPI
Each node in the supercomputer is assigned a unique node identity number (ID) ranging from 00 to $N-1$. When nodes are appropriately ordered, messages sent between nodes with closer IDs tend to require fewer network hops than those between nodes with larger ID differences.
@@
Leveraging this property, we sort the selected $\alpha$ bitstrings according to their excitation levels relative to the Hartree–Fock (HF) $\alpha$ bitstring, and assign them to nodes in increasing order. Hence, lower-excitation $\alpha$ bitstrings are allocated to lower-numbered nodes, while higher-excitation bitstrings are assigned to higher-numbered nodes.
This excitation-aware mapping can significantly reduce MPI communication between distant nodes. Since the excitation differences between the $\alpha$ bitstrings assigned to distant nodes are more likely to exceed two, remote data retrieval is usually unnecessary. As a result, and quite remarkably, even in massively parallel jobs involving thousands or tens of thousands of nodes, each node primarily fetches data from its neighboring nodes. This locality-preserving strategy plays a critical role in alleviating MPI communication congestion and enables our SCI program to maintain high parallel efficiency even on computations involving thousands or tens of thousands of nodes.
2.4.4 Achieving Memory and Computational Load Balance in MPI
As shown in Algorithm 1, the computational cost associated with each segment corresponding to a given $S^{\alpha}_{w}$ can be readily estimated. Consequently, the total workload of a process can be approximated by summing the estimated costs of all $\alpha$ bitstrings assigned to it. As described above, we assign $\alpha$ bitstrings to processes in increasing order of their excitation levels relative to the HF $\alpha$ bitstring, a strategy that is crucial in our parallel execution. Based on this excitation-aware mapping, achieving simultaneous balance in memory usage and computational workload becomes a nontrivial problem.
@@ 2.4.4 Achieving Memory and Computational Load Balance in MPI
Since the total number of $\alpha$ bitstrings is fixed, the average number per process is predetermined. Likewise, the total estimated cost is fixed, implying that the average computational load per process is also predetermined. Therefore, we can define two expansion factors to quantify the tradeoff between memory and computation: the memory expansion factor, defined as the ratio of the assigned $\alpha$ bitstring count to the average, and the computation expansion factor, defined as the ratio of the estimated computational cost to the average. Enforcing strict memory balance (i.e., keeping the memory expansion factor close to 1) can result in severe computational imbalance, causing some processes to complete significantly earlier and remain idle while others continue. Conversely, enforcing strict computational balance (i.e., keeping the computation expansion factor close to 1) may lead to a highly uneven memory usage, with some nodes exceeding their memory limits, thereby forcing an increase in the total number of nodes and leading to inefficient resource usage.
@@
To address this issue, we adopt a balanced strategy where both expansion factors are maintained to be moderately greater than 1, ensuring that neither memory imbalance nor computational imbalance becomes excessive. In the special case of FCI calculations—used here solely for testing purposes—the computational cost per segment is exactly the same, and thus a simple uniform distribution strategy naturally ensures both memory and workload balance.
2.4.5 Overlapping Computation and Data Transfer
@@
To mitigate this issue, one core is assigned to perform MPI data transfers, fetching the required data for the next step, while the remaining 47 cores perform parallel computation for the current step using OpenMP. Let $t_{\text{cal}}$ and $t_{\text{data}}$ represent the computation time and the data retrieval time of one step, respectively. If $t_{\text{data}}>t_{\text{cal}}$, we define the delay time to the wall time as $t_{\text{delay}}=t_{\text{data}}-t_{\text{cal}}$. Conversely, if $t_{\text{data}}\leq t_{\text{cal}}$, then $t_{\text{delay}}=0$, since MPI communication congestion during data transfer—if it occurs—does not prolong the wall time. The total delay time $T_{\text{delay}}$ for a Davidson iteration is defined as the sum of $t_{\text{delay}}$ over all $N$ steps of a process. As the wall time is determined by the slowest process, we only report the $T_{\text{delay}}$ observed on that process in the Section 3. As the single-precision implementation involves only half the data volume of the double-precision case, its $T_{\text{delay}}$ is expected to be significantly smaller, as confirmed by our numerical experiments.
2.4.6 Reassignment of $[0,2]$ for Delay Absorption
Since data retrieval and computation are executed concurrently, delays ($t_{\text{delay}}>0$) are more likely to occur in steps where the computational time $t_{\text{cal}}$ is relatively small. Through careful analysis, we observe that the $[0,2]$ term (Lines 27–32 in Algorithm 1) exhibits a favorable computational pattern: while being one of the dominant terms in both SCI and FCI calculations, it incurs substantial computational cost but only negligible data retrieval.
We therefore treat the $[0,2]$ tasks as a computational reservoir to absorb potential delays. Each process reassigns its $N$ batches of $[0,2]$ tasks to steps where $t_{\text{cal}}$ is pre-estimated to be small, thereby reducing both the probability and magnitude of $t_{\text{delay}}$. As discussed in Section 2.4.3, a local process may not need to retrieve any data from certain target processes due to the excitation-aware mapping. In such cases, delays are unlikely to occur, and the corresponding $[0,2]$ task for those steps are reassigned to other steps with small $t_{\text{cal}}$.
Due to this reassignment performed prior to execution, the actual execution no longer strictly follows the “one step computes one batch of DOUBLE_LINK” structure outlined in Algorithm 1. Instead, some steps may process multiple batches of $[0,2]$ tasks, some may process none, and some may be entirely skipped by process $p$. Nevertheless, for clarity, we still retain the original form of Algorithm 1.
2.4.7 Odd-Even Process Fetch Strategy
A large number of simultaneous MPI data transfers can lead to communication peaks in the network. To alleviate this issue, we introduce an odd–even process fetch strategy. For each process, the target ranks are sorted by the number of required segments. Odd-numbered processes fetch data in ascending order, while even-numbered processes fetch in descending order. This simple strategy effectively reduces both the frequency and magnitude of $T_{\text{delay}}$, particularly in simulations involving thousands or tens of thousands of nodes.
@@ 2.4.7 Odd-Even Process Fetch Strategy
2.4.9 Sleep Strategy under Severe MPI Congestion
Even with the above MPI optimization strategies, we still sometimes observe unexpectedly large values of $T_{\text{delay}}$ when using more than 5,000 nodes. Surprisingly, for the same system and input settings, $T_{\text{delay}}$ can be nearly zero in one run, but over 100 seconds in another. This instability is partly caused by external factors, such as MPI traffic from other jobs running at the same time, or different node placements by the job scheduler. However, we believe the deeper reason lies in the nature of MPI congestion. It is like a traffic jam: once congestion occurs, it induces blocking behavior, which in turn worsens further congestion. This kind of chain reaction becomes harder to control when using more nodes.
To address this, we design a sleep strategy. Once a process finds that $t_{\text{delay}}>0.2$ seconds for a given step, it is interpreted as an indication of severe congestion. The process then sleeps for 0.1 seconds to mitigate further blocking. While sleeping, the process is marked as BUSY, so that other processes will preferentially fetch data from other sources. Naturally, the sleep time is included in the measured $T_{\text{delay}}$.
Unlike the previous strategies, this sleep strategy may slightly increase the total $T_{\text{delay}}$ of a single Davidson iteration. However, in all of our tests, it consistently eliminates large delay events, making it a practical and effective safeguard. The specific thresholds of 0.2 and 0.1 seconds are chosen empirically based on our experience on Fugaku; different systems and applications may require fine-tuning of these parameters.
@@ 3 Results
We first generated the HF reference states and corresponding FCIDUMP files using PySCF. 22 Subsequently, TBSCI calculations were performed on the supercomputer Fugaku. The convergence criterion was defined as an energy difference of less than $2\times 10^{-7}$ Hartree between consecutive iterations. It should be noted that all the FCI data presented in this work were recomputed using our TBSCI program, although some of the results have been previously reported. 23, 24 Fugaku features a CPU-only architecture based on the A64FX processor. Each node contains 48 cores and only 32 GiB of memory, making both computational and memory distribution essential for large-scale execution. The TBSCI code was compiled using the Fujitsu compiler with the -Kfast,parallel,openmp,SVE flags. To demonstrate the performance of our code, we report the time of a single distributed matrix-vector multiplication ($\mathbf{W}=H\cdot\mathbf{U}$) for different numbers of processes; the wall time of this multiplication is almost as same as that of one Davidson iteration. Throughout all calculations, all data, including the CI vector and intermediate arrays, were kept entirely in memory without disk I/O, ensuring minimal data access latency.
@@ 3.1 N2 with the aug-cc-pVDZ Basis Set
We first test the performance of TBSCI by computing the potential energy surface (PES) of N2 using the aug-cc-pVDZ basis set. Under D2h symmetry and the frozen-core approximation, each geometry point yields $1.47\times 10^{11}$ determinants.
Table 2: Potential energy surface of N2 computed with the aug-cc-pVDZ basis set at various bond lengths ($R$). HF and FP64 (double-precision) FCI energies are reported in Hartree ($E_{h}$), and FP32 (single-precision) deviations from FP64 are shown in units of $10^{-6}\ E_{h}$.
$R$ (Bohr) HF FP64 FP32
2.068 -108.961045 -109.295263 1
2.4 -108.875323 -109.258928 1
2.7 -108.748038 -109.182160 0
3.0 -108.618929 -109.108790 0
3.6 -108.401280 -109.016333 -3
4.2 -108.242916 -109.986039 -9
[…]
@@
Table 3: Parallel performance of TBSCI for N2 computed with the aug-cc-pVDZ basis set using both double- and single-precision (FP64 and FP32) implementations. For each case, we report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.5) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
Nodes FP64 FP32
Avg. Max. $T_{\text{delay}}$ NHs Avg. Max. $T_{\text{delay}}$ NHs
480 2030 2088 0 278 1942 1997 0 266
960 1016 1053 0.1 281 968 1007 0 269
1440 676 704 0.3 282 645 673 0 269
1920 509 531 10.4 283 484 505 0 269
2400 407 426 4.2 284 387 405 0 270
2880 341 358 0.7 286 325 340 0 272
The MPI performance of both precision variants is summarized in Table 3. We begin by analyzing the MPI communication delay, characterized by the total $T_{\text{delay}}$ on the slowest process. Both precision variants show satisfactory performance in this regard. Remarkably, $T_{\text{delay}}$ is exactly zero in all single-precision runs, highlighting the benefit of reduced data volume in mitigating MPI congestion. Next, we assess MPI load balancing using the ratio of maximum to average wall time. Across all node counts, this ratio remains within 105%, confirming that the uniform distribution of $\alpha$ bitstrings achieves both memory and computational balance in FCI calculations. Finally, we evaluate strong scaling efficiency by tracking the total node hours (NHs) with increasing node counts. As shown in Table 3, NHs increase only slightly, indicating that the implementation maintains near-ideal strong scaling up to almost 3,000 nodes.
3.2 CN with the cc-pVTZ Basis Set
The next tested system is an open-shell molecule, CN. Under C2v symmetry with frozen core orbitals, the number of determinants reaches $2.46\times 10^{8}$ and $4.86\times 10^{11}$ under the cc-pVDZ and cc-pVTZ basis sets, respectively.
Table 4: Potential energy surfaces of CN computed with the cc-pVDZ and cc-pVTZ basis sets. HF and FP64 FCI energies are reported in Hartree ($E_{h}$), and FP32 deviations from FP64 are shown in units of $10^{-6}\ E_{h}$. The FP64 energies are extrapolated to the complete basis set (CBS) limit using a two-point formula, denoted as CBS($l$/$m$), where $l$ and $m$ denote the cardinalities of the basis sets.
$R$ (Å) cc-pVDZ cc-pVTZ CBS(2/3)
HF FP64 FP32 HF FP64 FP32 FP64
1.16 -92.197616 -92.491811 0 -92.219490 -92.569252 0 -92.595966
1.3 -92.149983 -92.479361 0 -92.168341 -92.550251 1 -92.575154
1.5 -92.041711 -92.408658 1 -92.059594 -92.476056 1 -92.499617
1.7 -91.937733 -92.334641 0 -91.957354 -92.400739 -1 -92.423284
2.0 -91.822293 -92.262054 -1 -91.844595 -92.324826 -1 -92.345249
2.3 -91.704156 -92.242169 4 -91.727232 -92.300497 -1 -92.318840
[…]
Table 5: Parallel performance of TBSCI for CN computed with the cc-pVTZ basis set using both FP64 and FP32 implementations. For each case, we report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
Nodes FP64 FP32
Avg. Max. $T_{\text{delay}}$ NHs Avg. Max. $T_{\text{delay}}$ NHs
1600 5516 5743 0 2552 5026 5265 0 2340
2400 3655 3809 0 2539 3331 3478 0 2319
4800 1824 1909 0.5 2545 1650 1743 0 2324
7200 1227 1315 18.1 2630 1110 1224 0 2448
9600 938 1070 8.3 2853 839 1003 0 2675
The MPI performance of TBSCI for CN with the cc-pVTZ basis set is summarized in Table 5. While the double-precision variant shows modest delays at high node counts (e.g., 18.1 seconds at 7200 nodes), the single-precision variant consistently shows zero delay on the slowest process in all tests—even at 9,600 nodes. Such excellent performance once again highlights the advantage of reduced data volume in large-scale high-performance computing (HPC) environments. At 9600 nodes, the ratio of maximum to average wall time increases to nearly 120%, indicating growing MPI load imbalance. Although in FCI calculations each segment should, in principle, have the same computational cost, such balance cannot be maintained once the workload is divided into multiple steps. Because perfect parallelism cannot be achieved within each step, increasing the number of steps—an inevitable consequence of using more nodes—leads to wall-time differences between processes. As a result, MPI load imbalance becomes more significant at larger node counts, leading to the observed increase in NHs. Nevertheless, from 1600 to 9600 nodes, the total NHs increase by only 14%, highlighting the exceptional performance enabled by the MPI optimization strategies we have designed.
3.3 Cr2 with the STO-3G Basis Set
Cr2 is a challenging molecule in quantum chemistry, exhibiting strong static and dynamic correlation even near its equilibrium geometry. With the minimal STO-3G basis set, the FCI dimension still reaches $9.14\times 10^{11}$ under D2h symmetry with 12 core orbitals frozen.
Table 6: Energies of Cr2 computed with various methods using the STO-3G basis set at $R=1.5$.
Methods Energy
HF -2064.078900
CCSD(T) -2064.781602
CCSDT -2064.780060
CCSDTQ -2064.802602
FP64 -2064.808195
FP32 -2064.808195
[…]
Table 7: Parallel performance of TBSCI for Cr2 computed with the STO-3G basis set using the single-precision implementation. We report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.5) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.